# A New Probe of Cosmic Birefringence Using Galaxy Polarization and Shapes

### v1

@@ Appendix D Radio luminosity function
where the factor $\lambda(z)$ accounts for the redshift evolution of spiral galaxies. For this, we adopt the simple formula from [63] that fits observed counts (also see [50]):
@@
The number density of spiral galaxies appear to drop quickly above $z=1.5$ [57], so we set $\phi(z,\log L)=0$ for $z\geq 1.5$.
@@
by exploiting the power-law frequency dependence of the spectral flux density with respect to frequency: $S_{\nu}\propto\nu^{\alpha}$, where the spectral index has an appropriate value $\alpha=-0.7$ [51, 52]. While [24] reported the empirical polarization-shape correlation at 4.84.84.8\,4.8GHz, we would like to relate sources at 4.84.84.8\,4.8GHz to sources at $\nu=1.4\,$GHz in order to make contact with observed counts.

### v2

@@ Appendix D Radio luminosity function
where the factor $\lambda(z)$ accounts for the redshift evolution of spiral galaxies. For this, we adopt the simple formula from Ref. [50] that fits observed counts:
@@
The number density of spiral galaxies appear to drop quickly above $z=1.5$ [61, 62], so we set $\phi(z,\log L)=0$ for $z\geq 1.5$.
@@
by exploiting the power-law frequency dependence of the spectral flux density with respect to frequency: $S_{\nu}\propto\nu^{\alpha}$, where the spectral index has an appropriate value $\alpha=-0.7$ [51, 52] (the readers should distinguish between this power-law index and the polarization rotation angle). While [24] reported the empirical polarization-shape correlation at 4.84.84.8\,4.8GHz, we would like to relate sources at 4.84.84.8\,4.8GHz to sources at $\nu=1.4\,$GHz in order to make contact with observed counts.